OpenStudy (anonymous):

Y = - 2/3x -6 Find the equation of the line that is perpendicular to this line and passes through the point (-3,5) . Find the equation of the line that is parallel to this line and passes through the point (-3,5)

OpenStudy (anonymous):

To find the equation of any linear function (equation of a line) you may use a point and a slope and use y=mx + b When you are told to find the equation of the line that is "perpendicular" to y=-2/3 x -6 you are being given a slope (indirectly) of that line. The slope of our given line is the number in front of x, so m= -2/3 The line that is perpendicular to y has slope -1/m (the negative reciprocal of m) so -1/m = -1/ (-2/3) = 3/2 so you want to find the equation of the line that has slope 3/2 and goes through the point (-3,5)

OpenStudy (anonymous):

so you have Y= Mx + b Y=3/2 x + b To find b you plug in the point (-3,5) 5= 3/2 (-3) + b 5= -9/2 +b 5 + 9/2 =b (10+9)/2=b 19/2=b Plug this into Y=3/2 x +b Y=3/2 x + 19/2

OpenStudy (anonymous):

For the second question, to find a line that is parallel that means we have the same slope...we want the equation of the line with slope -2/3 and that goes through (-3,5) Steps are similar to above :-)

OpenStudy (anonymous):

Thanks! It's been the better part of a decade since I've done any of this. I know it's hidden away in my brain someplace, just need some reminders.